OpenStudy (anonymous):

Given the function f(x) = x 2 + 4x - 3, find the value of f(a + 1). a2+ 4a- 3 2x2- 6x- 3 x2- 5x+ 3 a2+ 6a+ 2

OpenStudy (anonymous):

Do you want the answer by itself or with steps?

OpenStudy (anonymous):

\[\begin{split} & f(a+1) \\ =& (a+1)^2+4(a+1)-3 \\ =& (a+1)(a+1)+4(a+1)-3 \end{split}\]

OpenStudy (anonymous):

which one best for you

OpenStudy (anonymous):

No, don't give direct answer.

OpenStudy (anonymous):

It depends. It will be easy to give you the answer but I want you to understand.

OpenStudy (anonymous):

lol give the answer

OpenStudy (anonymous):

@wio Said I shouldn't I can't decide :P

OpenStudy (anonymous):

please give it to me

OpenStudy (anonymous):

f(a+1)=(a+1)2+4(a+1)−3 =(a+1)(a+1)+4(a+1)−3 You can take it from here :)

OpenStudy (anonymous):

f(a+1)=(a+1)^2+4(a+1)−3 =(a+1)(a+1)+4(a+1)−3 =(a^2 +2a +1) + 4a + 1 -3 K take it from here